OpenStudy (anonymous):

Find the center and vertices of the hyperbola 16x^2 - 9y^2 - 96x + 72y - 144 = 0

OpenStudy (anonymous):

your job is to make this look like \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] do you know how to do that?

OpenStudy (anonymous):

no I don't.. sorry i'm really lost!

OpenStudy (anonymous):

ok we can go slow, but the first thing to know is that if you want the center and vertices you need to put it in the form i wrote above, that way you can read off the answer if you make it look like \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\] then the center is \((h,k)\)

OpenStudy (anonymous):

this annoying task is done by "completing the square" do you know how to do that? if the answer is "no" that is fine i will walk you through it

OpenStudy (anonymous):

actually this one will not be so bad, because it was cooked up to give nice whole number answers ready?

OpenStudy (anonymous):

yeah I need help completing the square

OpenStudy (anonymous):

ok so the first thing to do it get the constant \(-144\) on the right side of the equal sign, and also put the \(x\) terms and \(y\) terms next to each other

OpenStudy (anonymous):

i.e. write \[16x^2-96x-9y^2+72y=144\]

OpenStudy (anonymous):

ok so far?

OpenStudy (anonymous):

Yep sounds good

OpenStudy (anonymous):

next step is to factor out the leading coefficient from the \(x\) terms and the \(y\) terms, which luckily is easy in this case

OpenStudy (anonymous):

\[16(x^2-6x)-9(y^2-8y)=144\]

OpenStudy (anonymous):

Oh okay i'm good so far

OpenStudy (anonymous):

notice that when you factor out the \(9\) from the \(y\) terms it is \(-9(y^2-8y)\) by the distributive law, not for example \(-9(y^2+8y)\) a common mistake

OpenStudy (anonymous):

now to complete the square have of 6 is 3, and half of 8 is 4 so write \[16(x-3)^2-9(y-4)^2=144+something\] we need to be careful about the something

OpenStudy (anonymous):

but before we find the 'something" is the left hand side clear? i mean clear how to get it? it is not that hard in this case

OpenStudy (anonymous):

Yeah I get it so far thank you

OpenStudy (anonymous):

ok what comes next doesn't usually happen, so let walk through it in any case we need to know what to add to the right, which is the same thing we added to the left when we completed the square when you change \(x^2-6x\) in to \((x-3)^2\) you have added 9, because \((x-3)^2=x^2-6x+9\) then when you multiply by 16 you see you have actually added \(9\times 16=144\) you don't have to do all this thinking, you just need to think \(3^2\times 16=144\) so if you add 144 to the left, you have to add 144 to the right

OpenStudy (anonymous):

then for the \(y\) term you have added \(4^2\times -9=-144\)

OpenStudy (anonymous):

so in fact in this example when you completed the square you added NOTHING to the elft and right because \(144-144=0\) but this is very UNUSUAL so don't think it always works this way, almost always you have to add something to the right

OpenStudy (anonymous):

now we are left with \[16(x-3)^2-9(y-4)^2=144\] to set it equal to 1, divide both sides by 144 and write \[\frac{(x-3)^2}{9}-\frac{(y-4)^2}{16}=1\]

OpenStudy (anonymous):

almost done you still following more or less?

OpenStudy (anonymous):

Yeah it makes sense so far!

OpenStudy (anonymous):

ok so we need the center, which is just a matter of reading\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]\[\frac{(x-3)^2}{9}-\frac{(y-4)^2}{16}=1\]so \(h=3,k=4\) and the center is \((3,4)\)

OpenStudy (anonymous):

Okay I get that thanks, how do you get the vertices?

OpenStudy (anonymous):

what else you need? foci or vertices?

OpenStudy (anonymous):

oh ok vertices

OpenStudy (anonymous):

for this we need to have some good idea what it looks like since \(9<16\) it look like this (roughly) 

OpenStudy (anonymous):

you need to know it look like this to see that the vertices are to the left and right of the center, not up and down

OpenStudy (anonymous):

in this example \(a^2=9\) and so \(a=3\) therefore you go 3 units to the left and right of the center the center is \((3,4)\) three units to the left is \((0,4)\) and 3 units to the right is \((6,4)\)

OpenStudy (anonymous):

hope it is clear that all the information comes from looking at \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]

OpenStudy (anonymous):

Yeah I understand that thank you so much

OpenStudy (anonymous):

yw